Copper has a mass density of 8.95g/cm3 and in electrical resistivity of 1.55 × 10−8ohm-m at room temperature. Assuming the effective mass is “mo”. Calculate.
1) The concentration of the conduction electrons
2) Electrical conductivity
1
Expert's answer
2020-06-25T09:36:05-0400

1)


"n_e=z\frac{\rho_m N_A}{\mu}=1\frac{(8950)(6.02\cdot 10^{26})}{63.55}=8.48\cdot 10^{28}\frac{el}{m^3}"

2)


"\sigma=\frac{1}{\rho }=\frac{1}{1.55\cdot 10^{-8}}=6.45\cdot 10^{7}\frac{1}{\Omega m}"